OpenStudy (anonymous):

What is the slope of a line that is perpendicular to the line determined by the equation 5x + 8y = 17?

TheSmartOne (thesmartone):

First make 5x+8y=17 into the format of y=mx+b 8y=-5x+17 y=-5/8x+17/83 Now the slopes of perpendiculars lines when multiplied toghether equals to -1 So \(\frac{-5}{8}\times x=-1\) Finally \(\ x=\frac{ 8 }{ 5 }\)